# Models for rational $(\infty,1)$-categories

### v1

@@ 8. Rational complete Segal spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Now, we introduce $M$-local complete Segal spaces, which also model $(\infty,1)$-categories enriched in $M$-local spaces. In particular, we introduce rational complete Segal spaces, which model rational $(\infty,1)$-categories. We left Bousfield localize Rezk’s model category for complete Segal spaces to obtain a model category whose fibrant objects are the $M$-local complete Segal spaces. Thus, we produce a model category whose fibrant objects are the rational complete Segal spaces. Then, we show that our model category for $M$-local complete Segal spaces remains compatible with the cartesian closure of simplicial spaces.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A complete Segal space $W$ is $M$-local if, for any two objects $x$ and $y$ of $W$, their mapping space $\operatorname{map}_{W}(x,y)$ is $M$-local. A complete Segal space $W$ is rational if, for any two objects $x$ and $y$ of $W$, their mapping space $\operatorname{map}_{W}(x,y)$ is rational.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $W$ be a complete Segal space. The following statements are equivalent.
@@
    The complete Segal space $W$ is $M$-local.
@@
    The Reedy fibrant Segal category $R(W)$ is $M$-local.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r},R(W))\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r},R(W))$ is a weak homotopy equivalence for every $i\geq 0$ and every $n\geq 2$.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r},W)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r},W)$ is a weak homotopy equivalence for every $i\geq 0$ and every $n\geq 2$.
@@
Thus, we localize Rezk’s model category for complete Segal spaces $\mathcal{CSS}$ at the same set
$\widetilde{T}_{M^{-1}\mathbb{Z}}=\{(\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r}\rightarrow(\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r}\mid i\geq 0\textrm{ and }n\geq 2\}.$
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a combinatorial, left proper, and simplicial model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$ on the category of simplicial spaces in which a map $f$ is
@@
    a weak equivalence if and only if, for every $M$-local complete Segal space $W$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(f,W)$ is a weak homotopy equivalence; and
@@
A simplicial space $W$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$ if and only if $W$ is an $M$-local complete Segal space.
@@
We apply the criterion from Proposition 5.5 to show that, for a multiplicative subset $M$ of $\mathbb{Z}$, our model category for $M$-local complete Segal spaces is cartesian. Our argument follows Rezk’s proof of [Rez01, Lemma 10.3], which is part of showing that the model category for complete Segal spaces is cartesian.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. If $W$ is an $M$-local complete Segal space, so is $W^{\Delta[1]^{t}}$.
@@
Firstly, Theorem 5.11 implies that $W^{\Delta[1]^{t}}$ is a complete Segal space. For $i\geq 0$ and $n\geq 2$, we show that
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r},W^{\Delta[1]^{t}})\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r},W^{\Delta[1]^{t}})$
is a weak homotopy equivalence. Equivalently, we show that the map
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[1]^{t}\times\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r},W)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[1]^{t}\times\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r},W)$
is a weak homotopy equivalence. For $0\leq j\leq i$, we define the order-preserving map $\gamma^{j}\colon[i+1]\rightarrow[1]\times[i]$ by
@@
For $0\leq j\leq i$, the inclusion maps
$(\gamma^{j}(\Delta[i+1]^{t})\times\operatorname{Sing}(S^{n}))_{r}\rightarrow(\gamma^{j}(\Delta[i+1]^{t})\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r}$
@@
$(\delta^{j}(\Delta[i]^{t})\times\operatorname{Sing}(S^{n}))_{r}\rightarrow(\delta^{j}(\Delta[i]^{t})\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r}$
are acyclic cofibrations in our model category for $M$-local complete Segal spaces, so the induced map of colimits
$(\Delta[1]^{t}\times\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r}\rightarrow(\Delta[1]^{t}\times\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r}$
is also an acyclic cofibration. As $W$ is an $M$-local complete Segal space, we conclude that the map
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[1]^{t}\times\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r},W)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[1]^{t}\times\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r},W)$
is a weak homotopy equivalence, as required. ∎
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. The model category for $M$-local complete Segal spaces $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$ is cartesian. In particular, the model category for rational complete Segal spaces $\mathcal{L}_{\mathbb{Q}}\mathcal{CSS}$ is cartesian.
@@ 9. Comparison of the two models for rational $(\infty,1)$-categories
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We show that the Quillen equivalence between complete Segal spaces and Segal categories descends to a Quillen equivalence between $M$-local complete Segal spaces and $M$-local Segal categories. Thus, we get a Quillen equivalence between rational complete Segal spaces and rational Segal categories, which captures that both structures are models for rational $(\infty,1)$-categories. Lastly, we prove that our model category for $M$-local Segal categories can be left-induced from that for $M$-local complete Segal spaces.
@@ 9.1. The equivalence of rational complete Segal spaces and rational Segal categories
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We show that our model category for $M$-local complete Segal spaces $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$ and our model category for $M$-local Segal categories $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ are Quillen equivalent. More specifically, our objective is to show that the Quillen equivalence $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{CSS}\colon R$ between Segal categories and complete Segal spaces from Theorem 6.9 descends to a Quillen equivalence $I\colon\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}\colon R$ between $M$-local Segal categories and $M$-local complete Segal spaces. To achieve such a descent, we need to left Bousfield localize both model categories in the Quillen equivalence from Theorem 6.9 in a compatible fashion as follows.
Theorem 9.1 ([Hir03, Theorem 3.3.20(1)(b)]).
@@
Remark 9.2.
If we relax the hypothesis that every object of $\mathcal{M}$ be cofibrant, then [Hir03, Theorem 3.3.20(1)(b)] reads that $F\colon\mathcal{L}_{T}\mathcal{M}\rightleftarrows\mathcal{L}_{\mathbb{L}F(T)}\mathcal{N}\colon G$ is a Quillen equivalence, where $\mathbb{L}F$ denotes the left derived functor of $F$; this functor involves applying $F$ to a suitable cofibrant replacement. Since the hypothesis that every object be cofibrant is satisfied by every model category in this paper, we add it to simplify the statement of the theorem.
As the left adjoint in the Quillen equivalence $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{CSS}\colon R$ is the inclusion and we left Bousfield localized both model categories at the same set, Theorem 9.1 yields our desired Quillen equivalence of localizations.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. The adjunction $I\colon\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}\colon R$ is a Quillen equivalence.
@@
The model category for rational complete Segal spaces and the model category for rational Segal categories are Quillen equivalent.
@@ 9.2. An approach to rational Segal categories by way of left transfer
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We conclude this paper by showing that, in analogy with our approach in Section 6.4, the adjunction from Theorem 9.3 can be used to left-induce our model category for $M$-local Segal categories $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ from our model category for $M$-local complete Segal spaces $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$. In particular, we show that we can left-induce our model category for rational Segal categories from our model category for rational complete Segal spaces.
Firstly, applying Theorem 6.11 to the adjunction $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}\rightleftarrows\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}\colon R$ yields the following model category.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a combinatorial model structure on the category of Segal precategories in which a map $\psi$ is
@@
    a weak equivalence if and only if $\psi$ is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$; and
@@
We verify the condition in Theorem 6.11: by our proof of Theorem 6.12, if a map $\psi$ of Segal precategories has the right lifting property against all monomorphisms of Segal precategories, then $\psi$ is a weak equivalence in $\mathcal{CSS}$, thus a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$. Now, Theorem 6.11 yields the model structure in our claim. ∎
It remains to verify that the model category from Proposition 9.5 is our model category for $M$-local Segal categories $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. In both model categories, the cofibrations are the monomorphisms, so it remains to prove that they also share the same weak equivalences. In other words, we need to show that a map of Segal precategories is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$ if and only if it is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. Both implications rest on the fact that we can replace a complete Segal space by a Segal category and vice versa up to Dwyer-Kan equivalence, and our arguments are underpinned by the following auxiliary lemma.
Lemma 9.6 ([Hir03, Corollary 9.3.3(1)]).
Let $\mathcal{M}$ be a simplicial model category. If $X$ is a cofibrant object and $f$ is a weak equivalence of fibrant objects, then $\operatorname{Map}_{\mathcal{M}}(X,f)$ is a weak homotopy equivalence.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. If a map of Segal precategories is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$, then it is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$.
@@
Let $\psi\colon X\rightarrow Y$ be a map of Segal precategories that is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. To prove that $\psi$ is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$, we show that, for every $M$-local complete Segal space $W$, the map
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(\psi,W)\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(Y,W)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(X,W)$
is a weak homotopy equivalence. Combining Lemma 9.6 with the Dwyer-Kan equivalence $R(W)\rightarrow W$ from Theorem 6.10, we equivalently show that, for every $M$-local complete Segal space $W$, the map
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(\psi,R(W))\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(Y,R(W))\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(X,R(W))$
is a weak homotopy equivalence, which is the case because $R(W)$ is a Reedy fibrant $M$-local Segal category by Proposition 8.2 and $\psi$ is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. Hence, $\psi$ is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$. ∎
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. If a map of Segal precategories is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$, then it is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$.
@@
Let $\psi\colon A\rightarrow B$ be a map of Segal precategories that is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$. We need to show that, for every Reedy fibrant $M$-local Segal category $X$, the map
@@
is a weak homotopy equivalence. As a Reedy fibrant Segal category is a Segal space, we apply Lemma 9.6 to the Dwyer-Kan equivalence $X\rightarrow\widehat{X}$ and show that, for every Reedy fibrant $M$-local Segal category $X$, the map
@@
is a weak homotopy equivalence, which is the case because $\widehat{X}$ is an $M$-local complete Segal space and $\psi$ is assumed to be a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$. Therefore, $\psi$ is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. ∎
Propositions 9.7 and 9.8 assemble to form the claim that we need to establish.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A map of Segal precategories is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{CSS}$ if and only if it is a weak equivalence in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$.
Proposition 9.9 is the $M$-local analog of Proposition 6.8, and it yields our promised left transfer.

### v2

@@ 7. Rational complete Segal spaces
Let $T$ be a set of simplicial maps. Now, we introduce $T$-local complete Segal spaces, which also model $(\infty,1)$-categories enriched in $T$-local spaces. In particular, we introduce rational complete Segal spaces, which model rational $(\infty,1)$-categories. We left Bousfield localize Rezk’s model category for complete Segal spaces to obtain a model category whose fibrant objects are the $T$-local complete Segal spaces. Thus, we produce a model category whose fibrant objects are the rational complete Segal spaces. Then, we show that our model category for $T$-local complete Segal spaces remains compatible with the cartesian closure of simplicial spaces.
@@
Let $T$ be a set of simplicial maps. A complete Segal space $W$ is $T$-local if, for any two objects $x$ and $y$ of $W$, their mapping space $\operatorname{map}_{W}(x,y)$ is $T$-local. A complete Segal space $W$ is rational if, for any two objects $x$ and $y$ of $W$, their mapping space $\operatorname{map}_{W}(x,y)$ is rational.
@@
Let $T$ be a set of simplicial maps. Let $W$ be a complete Segal space. The following statements are equivalent.
@@
    The complete Segal space $W$ is $T$-local.
@@
    The Reedy fibrant Segal category $R(W)$ is $T$-local.
@@
    For every $i\geq 0$ and every $\lambda\in T$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\lambda)_{r},R(W))$ is a weak equivalence.
@@
    For every $i\geq 0$ and every $\lambda\in T$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\lambda)_{r},W)$ is a weak equivalence.
@@
Thus, we localize Rezk’s model category for complete Segal spaces $\mathcal{CSS}$ at the set $\widetilde{T}$ from Section 6.
@@
Let $T$ be a set of simplicial maps. There is a combinatorial, left proper, and simplicial model structure $\mathcal{L}_{T}\mathcal{CSS}$ on the category of simplicial spaces in which a map $f$ is
@@
    a weak equivalence if and only if, for every $T$-local complete Segal space $W$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(f,W)$ is a weak equivalence; and
@@
A simplicial space $W$ is fibrant in $\mathcal{L}_{T}\mathcal{CSS}$ if and only if $W$ is a $T$-local complete Segal space.
@@
We apply the criterion from Proposition 4.5 to show that, for a set of simplicial maps $T$, our model category for $T$-local complete Segal spaces is cartesian. Our argument follows Rezk’s proof of [25, Lemma 10.3], which is part of showing that the model category for complete Segal spaces is cartesian.
@@
Let $T$ be a set of simplicial maps. If $W$ is a $T$-local complete Segal space, so is $W^{\Delta[1]^{t}}$.
@@
Replacing maps in $T$ by monomorphisms up to weak equivalence if necessary, we assume without loss of generality that $T$ comprises monomorphisms. By Theorem 4.11, $W^{\Delta[1]^{t}}$ is a complete Segal space. For $i\geq 0$ and a map $A\rightarrow B$ in $T$, we show that
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times B)_{r},W^{\Delta[1]^{t}})\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times A)_{r},W^{\Delta[1]^{t}})$
is a weak equivalence. Equivalently, we show that the map
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[1]^{t}\times\Delta[i]^{t}\times B)_{r},W)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[1]^{t}\times\Delta[i]^{t}\times A)_{r},W)$
is a weak equivalence. For $0\leq j\leq i$, we define the order-preserving map $\gamma^{j}\colon[i+1]\rightarrow[1]\times[i]$ by
@@
For $0\leq j\leq i$, because the set $T$ comprises monomorphisms, the inclusion maps
$(\gamma^{j}(\Delta[i+1]^{t})\times A)_{r}\rightarrow(\gamma^{j}(\Delta[i+1]^{t})\times B)_{r}$
@@
$(\delta^{j}(\Delta[i]^{t})\times A)_{r}\rightarrow(\delta^{j}(\Delta[i]^{t})\times B)_{r}$
are acyclic cofibrations in our model category for $T$-local complete Segal spaces, so the induced map of colimits
$(\Delta[1]^{t}\times\Delta[i]^{t}\times A)_{r}\rightarrow(\Delta[1]^{t}\times\Delta[i]^{t}\times B)_{r}$
is also an acyclic cofibration. As $W$ is a $T$-local complete Segal space, our claim follows. ∎
Remark 7.6.
To relax the hypothesis that the set $T$ comprise monomorphisms, we would ultimately need the class of weak equivalences in a model category to be closed under colimits, which is not the case; a counterexample using pushouts of spaces is presented at the start of [13, §10] to motivate the use of homotopy pushouts.
@@
Let $T$ be a set of simplicial maps. The model category for $T$-local complete Segal spaces $\mathcal{L}_{T}\mathcal{CSS}$ is cartesian. In particular, the model category for rational complete Segal spaces is cartesian.
@@ 8. Comparison of the two models for rational $(\infty,1)$-categories
Let $T$ be a set of simplicial maps. We show that the Quillen equivalence between complete Segal spaces and Segal categories descends to one between $T$-local complete Segal spaces and $T$-local Segal categories. Thus, we get a Quillen equivalence between rational complete Segal spaces and rational Segal categories, which captures that both structures are models for rational $(\infty,1)$-categories. Lastly, we prove that our model category for Reedy fibrant $T$-local Segal categories can be left-induced from that for $T$-local complete Segal spaces.
@@ 8.1. The equivalence of rational complete Segal spaces and rational Segal categories
Let $T$ be a set of simplicial maps. We show that our model category for $T$-local complete Segal spaces $\mathcal{L}_{T}\mathcal{CSS}$ and our model category for Reedy fibrant $T$-local Segal categories $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ are Quillen equivalent. Our objective is to show that the Quillen equivalence $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{CSS}\colon R$ between Segal categories and complete Segal spaces from Theorem 5.10 descends to a Quillen equivalence $I\colon\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{L}_{T}\mathcal{CSS}\colon R$ between $T$-local Segal categories and $T$-local complete Segal spaces. To achieve such a descent, we need to left Bousfield localize both model categories in the Quillen equivalence from Theorem 5.10 in a compatible fashion as follows.
Theorem 8.1 ([18, Theorem 3.3.20(1)(b)]).
@@
Observe that every object of $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ is cofibrant. As the left adjoint in $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{CSS}\colon R$ is the inclusion and we localized both model categories at the same set, Theorem 8.1 yields our desired Quillen equivalence.
@@
Let $T$ be a set of simplicial maps. The adjunction $I\colon\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightleftarrows\mathcal{L}_{T}\mathcal{CSS}\colon R$ is a Quillen equivalence.
@@
The model category for rational complete Segal spaces and the model category for Reedy fibrant rational Segal categories are Quillen equivalent.
@@ 8.2. An approach to rational Segal categories by way of left transfer
Let $T$ be a set of simplicial maps. We show that, in analogy with our approach in Section 5.4, the adjunction from Theorem 8.2 can be used to left-induce our model category for Reedy fibrant $T$-local Segal categories $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ from our model category for $T$-local complete Segal spaces $\mathcal{L}_{T}\mathcal{CSS}$. In particular, we left-induce our model category for Reedy fibrant rational Segal categories from our model category for rational complete Segal spaces.
Firstly, applying Theorem 5.12 to the adjunction $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}\rightleftarrows\mathcal{L}_{T}\mathcal{CSS}\colon R$ yields the following model category.
@@
Let $T$ be a set of simplicial maps. There is a combinatorial model structure on the category of Segal precategories in which a map $\psi$ is
@@
    a weak equivalence if and only if $\psi$ is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$; and
@@
We verify the condition in Theorem 5.12. By our proof of Theorem 5.13, if a map $\psi$ of Segal precategories has the right lifting property against all monomorphisms of Segal precategories, then $\psi$ is a weak equivalence in $\mathcal{CSS}$, thus a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$. Now, Theorem 5.12 yields the model structure in our claim. ∎
It remains to verify that the model category from Proposition 8.4 is our model category for Reedy fibrant $T$-local Segal categories $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. In both model categories, the cofibrations are the monomorphisms, so it remains to prove that they also share the same weak equivalences. In other words, we need to show that a map of Segal precategories is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$ if and only if it is a weak equivalence in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. Both implications rest on the fact that we can replace a complete Segal space by a Segal category and vice versa up to Dwyer-Kan equivalence. Our arguments are underpinned by the following lemma.
Lemma 8.5 ([18, Corollary 9.3.3(1)]).
Let $\mathcal{M}$ be a simplicial model category. If $X$ is a cofibrant object and $f$ is a weak equivalence of fibrant objects, then $\operatorname{Map}_{\mathcal{M}}(X,f)$ is a weak equivalence.
@@
Let $T$ be a set of simplicial maps. If a map of Segal precategories is a weak equivalence in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$, then it is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$.
@@
Let $\psi\colon A\rightarrow B$ be a map of Segal precategories that is a weak equivalence in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. To prove that $\psi$ is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$, we show that, for every $T$-local complete Segal space $W$, the map
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(\psi,W)\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(B,W)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(A,W)$
is a weak equivalence. Combining Lemma 8.5 with the Dwyer-Kan equivalence $R(W)\rightarrow W$ from Theorem 5.11, we equivalently show that, for every $T$-local complete Segal space $W$, the map
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(\psi,R(W))\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(B,R(W))\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(A,R(W))$
is a weak equivalence, which is the case because $R(W)$ is a Reedy fibrant $T$-local Segal category by Proposition 7.2 and $\psi$ is a weak equivalence in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. Hence, $\psi$ is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$. ∎
@@
Let $T$ be a set of simplicial maps. If a map of Segal precategories is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$, then it is a weak equivalence in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$.
@@
Let $\psi\colon A\rightarrow B$ be a map of Segal precategories that is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$. We need to show that, for every Reedy fibrant $T$-local Segal category $X$, the map
@@
is a weak equivalence. As a Reedy fibrant Segal category is a Segal space, we apply Lemma 8.5 to the Dwyer-Kan equivalence $X\rightarrow\widehat{X}$ and show that, for every Reedy fibrant $T$-local Segal category $X$, the map
@@
is a weak equivalence, which is the case because $\widehat{X}$ is a $T$-local complete Segal space and $\psi$ is assumed to be a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$. Therefore, $\psi$ is a weak equivalence in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. ∎
Propositions 8.6 and 8.7 assemble to form the claim that we need to establish.
@@
Let $T$ be a set of simplicial maps. A map of Segal precategories is a weak equivalence in $\mathcal{L}_{T}\mathcal{CSS}$ if and only if it is a weak equivalence in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$.
Proposition 8.8 yields our promised left transfer.